PLEASE HELP: I have to find the total amount due for each simple interest loan for $3125 at 4.81% interest for 10 years.: PLEASE HELP: I have to find the total amount due for each simple interest loan for $3125 at 4.81% interest for 10 years.
Answer by mangopeeler07(442)

First figure out what 4.81% of 3125 is. So set it up as
x/3125=4.81/100

Cross multiply
100x=15031.25

Divide by 100
x=150.3125

In money terms, that is
$150.31

Add that to $3125.00 ten times
3125+10(150.31)
3125+1503.10=4628.10

Amount due: $4628.10